The terminal side of an angle passes through the given ordered pairs below. For each angle, find the values of the six trigonometric ratios.

(2,-sqrt( 5 ))
sin⁡θ = ?
cosθ = ?
tanθ = ?
cscθ = ?
secθ = ?
cotθ = ?
Thanks!!

Answer by Alan3354(69443)

The terminal side of an angle passes through the given ordered pairs below. For each angle, find the values of the six trigonometric ratios.
-----------
x = 2
y = -sqrt(5)
r = sqrt(x^2 + y^2)
-----
sin = y/r
cos = x/r
tan = sin/cos
---
cot = 1/tan
sec = 1/cos
csc = 1/sin
========================
PS These are things you should know.